# Cluster algebras and quantum cohomology rings: A-type

### v1

@@ 4 Proof for all-genus Seiberg duality Conjecture
We will utilize the localization theorem to prove that all-genus Seiberg duality Conjecture 3.3 holds for all decorated QPs $(\mathbf{Q},\mathbf{r})\in\Omega^{A}_{n}$ that are mutation equivalent to $A_{n}$. Throughout this section, we will let $(\mathbf{Q},\mathbf{r})$ and $(\mathbf{Q}^{\prime},\mathbf{r}^{\prime})$ be two decorated QPs that are related by a quiver mutation $\mu_{v}$ in $\Omega^{A}_{n}$ and let $\mathcal{Z}$ and $\mathcal{Z}^{\prime}$ be the varieties of the two QPs.
@@
$\langle\tau_{a_{1}}\gamma_{1},\ldots,\tau_{a_{m}}\gamma_{m}\rangle_{g,m,\beta}^{\mathcal{Z},S}=\langle\tau_{a_{1}}\phi(\gamma_{1}),\ldots,\tau_{a_{m}}\phi(\gamma_{m})\rangle_{g,m,\phi_{*}(\beta)}^{\mathcal{Z}^{\prime},S}.$ (4.1)
See Proposition 4.9 and Proposition 4.10 for the constructions of the map $\phi$ and $\phi_{*}$.
@@
$q_{u}^{\prime}=q_{v}\prod_{u\in Q_{0}\backslash Q_{f}}(q_{u})^{[\operatorname{sign}(\sigma_{v})b_{vu}]_{+}}\,.$ (4.2)
@@ 4.1 Localization theorem
Localization is a very efficient tool to compute Gromov-Witten invariants of a GKM space, and we review it in this section following [GP99, Liu11]. A variety $\mathcal{X}$ with a torus action $S$ is a GKM space if the set of zero-dimensional $S$-orbits is of zero dimension and the set of one-dimensional $S$-orbits is of one dimension. Under this assumption, the torus fixed locus $\mathcal{X}^{S}$ is discrete and the closure of any irreducible 1-dimensional $S$-orbit in $\mathcal{X}$ is isomorphic to $\mathbb{P}^{1}$, with 00 and $\infty$ being $S$-fixed.
We define the 1-skeleton of $\mathcal{X}$ to be the union of all closures of one-dimensional orbits, which we denote by $\mathcal{X}^{1}$.
@@ 4.1.1 Torus fixed graphs
We assign a graph (a compact, connected 1-dimensional CW complex) $\Upsilon$ to the 1-skeleton $\mathcal{X}^{1}$. The vertices $\sigma$ in $\Upsilon$ are $S$-fixed points $p_{\sigma}\in\mathcal{X}^{S}$, and edges ${\tau}$ in $\Upsilon$ are irreducible 1-dimensional $S$-orbits $\ell_{\tau}$. A vertex $\sigma$ is contained in an edge $\tau$ if and only if the fixed point $p_{\sigma}$ is contained in the closure of the 1-dimensional $S$-orbit $\ell_{\tau}$. A flag in $\Upsilon$ is defined as a pair $(\tau,\sigma)\in\mathcal{X}^{S}\times\mathcal{X}^{1}$ where the vertex $\sigma$ is contained in the edge $\tau$. We denote the sets of vertices, edges, and flags of $\Upsilon$ by $V(\Upsilon)$, $E(\Upsilon)$ and $F(\Upsilon)$.
@@
$\mathbf{w}(\tau,\sigma)=c_{1}^{S}(T_{p_{\sigma}}\ell_{\tau})\in H^{2}_{S}(pt).$ (4.3)
@@
    $\mathbf{w}(\tau_{i}`,\sigma`)=\mathbf{w}(\tau_{i},\sigma)-a_{i}\mathbf{w}(\tau,\sigma).$ (4.4)
@@ 4.1.2 Decorated graphs
In this section, we deal with torus fixed points in the moduli stack of stable maps ¯$\overline{\mathcal{M}}_{g,n}(\mathcal{X},\beta)$. Given a stable map $f:(C,x_{1},\cdots,x_{n})\rightarrow\mathcal{X}$ such that $[f:(C,x_{1},\cdots,x_{n})\rightarrow\mathcal{X}]\in\overline{\mathcal{M}}_{g,n}(\mathcal{X},\beta)^{S}$, we will associate to it a decorated graph $\overrightarrow{\Gamma}$. We first give a formal definition.
@@
(i) $\sum_{v\in V(\Gamma)}g_{v}+|E(\Gamma)|-|V(\Gamma)|+1=g$.
(ii) $\sum_{e\in E(\Gamma)}d_{e}[\ell_{\tau_{e}}]=\beta$.
Let $G_{g,n}(\mathcal{X},\beta)$ be the set of all decorated graphs $\overrightarrow{\Gamma}=(\Gamma,\overrightarrow{f},\overrightarrow{d},\overrightarrow{g},\overrightarrow{s})$ satisfying the above constraints.
@@
the set of edges emanating from $v$, and define $S_{v}=\overrightarrow{s}^{-1}(v)\subset\{1,\ldots,n\}$. The valency of $v$ is defined by $val(v)=|E_{v}|+|S_{v}|$.
Let $f:(C,x_{1},\cdots,x_{n})\in\overline{\mathcal{M}}_{g,n}(\mathcal{X},\beta)^{S}$ be a stable map fixed by $S$-action. Then for an irreducible component $C^{\prime}\subset C$, $f\rvert_{C^{\prime}}$ either contracts to a point $p_{\sigma}\in\mathcal{X}^{S}$ or $C^{\prime}\cong\mathbb{P}^{1}$ and $f(C^{\prime})=\ell_{\tau}$.
A decorated graph $\overrightarrow{\Gamma}$ associated to $f:(C,x_{1},\ldots,x_{n})\rightarrow\mathcal{X}$ can be defined as follows.
  1. (1)
    We assign a vertex $v$ to each connected component of $f^{-1}(\mathcal{X}^{S})$, and denote the corresponding component by $C_{v}$. Assume that $f(C_{v})=\{p_{\sigma}\}$, then the label map $\overrightarrow{f}(v)=\sigma$, and the genus $\overrightarrow{g}(v)=g_{v}$ is the arithmetic genus of the component $C_{v}$. The marking map $\overrightarrow{s}$ sends marking points on component $C_{v}$ to $v\in V(\Gamma)$.
  2. (2)
    We assign an edge $e$ to each connected component $f^{-1}(\mathcal{X}^{1}\backslash\mathcal{X}^{S})$. Let $C_{e}\cong\mathbb{P}^{1}$ be the closure of one connected component in $f^{-1}(\mathcal{X}^{1}\backslash\mathcal{X}^{S})$ and assume $f(C_{e})=\ell_{\tau}$. We define $\overrightarrow{f}(e)=\tau$.
  3. (3)
@@
The above construction gives a map from ¯$\overline{\mathcal{M}}_{g,n}(\mathcal{X},\beta)^{S}$ to the discrete set $\overrightarrow{\Gamma}\in G_{g,n}(\mathcal{X},\beta)$. Let $\mathcal{F}_{\overrightarrow{\Gamma}}\subset\overline{\mathcal{M}}_{g,n}(\mathcal{X},\beta)^{S}$ denote the preimage of $\overrightarrow{\Gamma}$. Then
@@ 4.1.3 Localization formula on Gromov-Witten theory
where $\Lambda^{\vee}_{g}(u)=\Sigma_{i=0}^{g}(-1)^{i}\lambda_{i}u^{g-i}$ and $\lambda_{i}$ are $\lambda$-classes, see the definition in [Liu11, Sec 3.1]. Furthermore, we define
@@
4.2 Torus fixed points of varieties of QPs in $\Omega^{A}_{n}$
@@
$(Fl)^{S}=\{\vec{I}=(I_{1},\cdots,I_{n})|I_{1}\subset I_{2}\subset\cdots\subset[N_{n+1}],\lvert I_{i}\rvert=N_{i},i=1,\cdots,n\}\,,$ (4.5)
@@
$\lvert(Fl)^{S}\rvert=\prod_{i=1}^{n}\binom{N_{i+1}}{N_{i}}.$ (4.6)
@@
  • If a node $v$ has only one maximal node $w$ smaller than it, the corresponding matrix is non-degenerate and in reduced row/column echelon form whose non-pivots entries vanish.
@@
Combine the Equation (4.6) and the Lemma 4.7, and we obtain the following Corollary.
@@
For an arbitrary decorated QP $(\mathbf{Q}=(Q_{0},Q_{1},W),\mathbf{r})\in\Omega^{A}_{n}$, let $\mathcal{Z}$ be the critical locus of $W$. The cardinality of the torus fixed locus satisfies
$\lvert\mathcal{Z}^{S}\rvert=\lvert(Fl)^{S}\rvert=\prod_{i=1}^{n}\binom{N_{i+1}}{N_{i}}$
According to the Atiyah-Bott localization theorem [AB84],
$H_{S}^{*}(\mathcal{Z})\cong\oplus_{p\in\mathcal{Z}^{S}}H^{*}_{S}(p)\,,$ (4.7)
@@
$\phi:H^{*}_{S}(\mathcal{Z})\xrightarrow{\sim}H^{*}_{S}(\mathcal{Z}^{\prime})\,.$ (4.8)
@@
$\gamma\big{|}_{P}=\phi(\gamma)\big{|}_{\varphi(P)},$ (4.9)
@@
It is obvious that the two equivariant cohomology groups are isomorphic by Atiyah-Bott localization theorem (4.7) and the Corollary 4.8. One can find the mutation $\mu_{v}$ maps $c_{k}^{S}(S_{u}^{\vee})$ to $c_{k}^{S}(S_{u}^{\prime})$ if $\sigma_{u}$ changes its sign according to the Definition 3.4.
@@
$\int_{\beta_{u}}c_{1}(S_{w})=-\operatorname{sign}(\sigma_{w})\delta_{uw}\,.$ (4.11)
Then we have $H_{2}(\mathcal{Z},\mathbb{Z})\cong\mathbb{Z}^{n}$ For two varieties $\mathcal{Z}$ and $\mathcal{Z}^{\prime}$ related by a quiver mutation at the gauge node $v$, we present a transformation of their effective curve classes inspired by (3.9).
@@
Let $\beta=(\beta_{u})_{u\in Q_{0}\backslash Q_{f}}$ and $\beta^{\prime}=(\beta^{\prime}_{u})_{u\in Q_{0}\backslash Q_{f}}$ be the bases of $H_{2}(\mathcal{Z},\mathbb{Z})$ and $H_{2}(\mathcal{Z}^{\prime},\mathbb{Z})$, respectively, chosen as (4.11). Then their transformation is given as follows. If $\sigma_{v}>0$ which means $N_{f}(v)>N_{a}(v)$, then
$\phi_{*}(\beta_{u})=\begin{cases}-\beta_{u}^{\prime},&\text{ for }u=v\\ \beta_{u}^{\prime}+\beta_{v}^{\prime}&\text{ for }v\rightarrow u\\ \beta_{u}^{\prime}&\text{ otherwise }\,.\end{cases}$ (4.12)
@@
$\phi_{*}(\beta_{u})=\begin{cases}-\beta_{u}^{\prime},&\text{ for }u=v\\ \beta_{u}^{\prime}+\beta_{v}^{\prime}&\text{ for }u\rightarrow v\\ \beta_{u}^{\prime}&\text{ otherwise }\,.\end{cases}$ (4.13)
@@
Notice that $(c_{1}(S_{u}))_{u\in Q_{0}\backslash Q_{f}}$ spans $H^{2}(\mathcal{Z})$. One can check directly that the $\phi_{*}$ defined above satisfies (3.9), via Proposition 4.9 and (4.11). ∎
4.3 One-dimensional $S$-orbits in varieties of QPs in $\Omega^{A}_{n}$
@@ 4.3.1 Grassmannian
    $A=\begin{bmatrix}\vec{v}_{1}&\vec{v}_{2}&\cdots&\vec{v}_{N_{2}}\end{bmatrix}.$ (4.14)
@@
    $\begin{cases}(A_{t})_{ab}=\delta_{f_{a}b}\text{ for }a\neq s\\ (A_{t})_{si_{s}}=\frac{1}{1+t}\\ (A_{t})_{sj_{s}}=\frac{t}{t+1}\end{cases}$ (4.15)
@@
${Fl(N_{1},\ldots,N_{n})}$${Fl(N_{1},\ldots,N_{n},N_{n+1})}$${Gr(N_{n},N_{n+1})}$$\scriptstyle{\pi}$ (4.16)
@@
  2. (2)
    If $\pi(\ell)$ is a line $\ell_{b}\subset Gr(N_{n},N_{n+1})$, then $\ell_{b}$ is an one-dimensional $S$-orbit closure in $Gr(N_{n},N_{n+1})$ since $\pi$ is $S$-equivariant. By Lemma 4.11, $\ell_{b}$ contains two $S$-fixed points $q_{I},q_{J}\in Gr(N_{n},N_{n+1})^{S}$, satisfying $|I\cap J|=N_{n}-1$. Without loss of generality, we assume that $I\backslash J=\{N_{n+1}-1\}$, $J\backslash I=\{N_{n+1}\}$. Fix a point $q\in\ell_{b}\backslash\{q_{I},q_{J}\}$, and we assume that its matrix representation is given by $A_{t_{0}}$ in (4.15) for some fixed $t_{0}$. The matrix representation of the fiber $\pi^{-1}(q)$ can be $(A_{1},A_{2},\ldots,A_{n-1},A_{t_{0}})$. Now one can check that there is a $S^{\prime}=(\mathbb{C}^{*})^{N_{n+1}-1}$-action on $\pi^{-1}(q)$, where the inclusion $S^{\prime}\hookrightarrow S$ is given by $(t_{1},t_{2},\ldots,t_{N-2},t,t)$. Actually, this action can be viewed as an extension of the natural $(\mathbb{C}^{*})^{N_{n}}$-action on $Fl(N_{1},N_{2},\ldots,N_{n})$ to $\pi^{-1}(q)$.
    The point $p\in\ell\cap\pi^{-1}(q)$ must be a $S^{\prime}$-fixed point in $Fl(N_{1},N_{2},\ldots,N_{n})$. Otherwise the whole $S$-action on $p$ will result in a larger than one dimensional $S$-orbit, which contradicts the assumption that $\ell$ is one-dimensional. By continuity, when $q$ varies, $p$ should have the same matrix representation. In the formula of (4.5), either $I_{l}=J_{l}$, or $I_{l}$ contains the $N_{n+1}-1$ and $J_{l}$ contains the $N_{n+1}$. Hence we have proved the induction in case $n$.
@@
    $A_{1}=\begin{bmatrix}s&t&0&\cdots&0&\ldots&0\\ 0&0&1&\cdots&0&\ldots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots&&\vdots\\ 0&0&0&\cdots&1&\ldots&0\end{bmatrix}$ (4.17)
@@
    $A_{2}=\begin{bmatrix}a&0&0&\cdots&0\\ b&0&0&\cdots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ 0&0&1&\cdots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ 0&0&0&\cdots&1\end{bmatrix}.$ (4.18)
@@
We consider a general decorated QP in $\Omega^{A}_{n}$. Let $\mathcal{Z}:=\{dW=0\}\sslash_{\theta}G$ be the associated variety. We have given the $S$-fixed locus $\mathcal{Z}^{S}$ in Lemma 4.6 indexed by a set of sets $\vec{I}=(I_{v})_{v\in Q_{0}\backslash v_{f}}$. Denote the corresponding $S$-fixed point by $p_{\vec{I}}$.
@@
  2. (2)
    There exist at most two maximal nodes that are smaller than $v_{0}$.
    If there is only one maximal node $v_{1}$ that is smaller than $v_{0}$, then $I_{v_{1}}\cap J_{v_{1}}=r_{v_{1}}-1$.
    If there are exactly two maximal nodes $v_{1}$ and $v_{2}$ that are smaller than $v_{0}$, then $(I_{v_{k}},J_{v_{k}}),\,k=1,2$ must satisfy one of the following three conditions:
@@
We prove by induction on the number of gauge nodes $n$.
For $n=1$ and $2$, this is proved by Lemmas 4.11, 4.12 and 4.13.
Assume that the Lemma is true for $n-1$.
We’re going to prove the case when the number of gauge node is $n$. Let $\ell$ be an one-dimensional $S$-orbit closure in $\mathcal{Z}$. When the frozen node $v_{f}$ has only one maximal node $v_{1}$ smaller than it, then we have the following fiber bundle
${\mathcal{Z}^{\prime}}$${\mathcal{Z}}$${Gr(r_{v_{1}},N_{n+1})}$$\scriptstyle{\pi}$ (4.19)
@@
4.4 Behavior of one-dimensional $S$-orbits under quiver mutations

### v2

@@ 4 Proof for all-genus Seiberg duality Conjecture
We will utilize the localization theorem to prove that the Seiberg duality conjecture of all genus 3.3 holds for all decorated QPs $(\mathbf{Q},\mathbf{r})\in\Omega^{A}_{n}$ that are mutation equivalent to $A_{n}$. Throughout this section, we will let $(\mathbf{Q},\mathbf{r})$ and $(\mathbf{Q}^{\prime},\mathbf{r}^{\prime})$ be two decorated QPs that are related by a quiver mutation $\mu_{v}$ in $\Omega^{A}_{n}$ and let $\mathcal{Z}$ and $\mathcal{Z}^{\prime}$ be the varieties of the two QPs.
@@
$\langle\tau_{a_{1}}\gamma_{1},\ldots,\tau_{a_{m}}\gamma_{m}\rangle_{g,m,\beta}^{\mathcal{Z},S}=\langle\tau_{a_{1}}\phi(\gamma_{1}),\ldots,\tau_{a_{m}}\phi(\gamma_{m})\rangle_{g,m,\phi_{*}(\beta)}^{\mathcal{Z}^{\prime},S}.$
See Proposition 4.9 and Proposition 4.10 for the constructions of maps $\phi$ and $\phi_{*}$.
@@
$q_{u}^{\prime}=q_{v}\prod_{u\in Q_{0}\backslash Q_{f}}(q_{u})^{[\operatorname{sign}(\sigma_{v})b_{vu}]_{+}}\,.$
@@ 4.1 Localization theorem
Localization is a very efficient tool for computing Gromov-Witten invariants of a GKM space, and we review it in this section following [GW:GPequiv, liu2011localization]. A variety $\mathcal{X}$ with torus action $S$ is a GKM space if the set of zero-dimensional $S$-orbits is of zero dimension and the set of one-dimensional $S$-orbits is of one dimension. Under this assumption, the torus fixed locus $\mathcal{X}^{S}$ is discrete and the closure of any irreducible 1-dimensional $S$-orbit in $\mathcal{X}$ is isomorphic to $\mathbb{P}^{1}$, with 00 and $\infty$ being $S$-fixed.
We define the 1 skeleton of $\mathcal{X}$ as the union of all closures of one-dimensional orbits, which we denote by $\mathcal{X}^{1}$.
@@ 4.1.1 Torus fixed graphs
We assign a graph (a compact, connected 1-dimensional CW complex) $\Upsilon$ to the 1-skeleton $\mathcal{X}^{1}$. The vertices $\sigma$ in $\Upsilon$ are $S$-fixed points $p_{\sigma}\in\mathcal{X}^{S}$, and edges ${\tau}$ in $\Upsilon$ are irreducible 1-dimensional $S$-orbits $\ell_{\tau}$. A vertex $\sigma$ is contained in an edge $\tau$ if and only if the fixed point $p_{\sigma}$ is contained in the closure of the 1-dimensional $S$-orbit $\ell_{\tau}$. A flag in $\Upsilon$ is defined as a pair $(\tau,\sigma)\in\mathcal{X}^{S}\times\mathcal{X}^{1}$ where the vertex $\sigma$ is contained in the edge $\tau$. We denote the sets of vertices, edges, and flags of $\Upsilon$ by $V(\Upsilon)$, $E(\Upsilon)$, and $F(\Upsilon)$.
@@
(12) $\mathbf{w}(\tau,\sigma)=c_{1}^{S}(T_{p_{\sigma}}\ell_{\tau})\in H^{2}_{S}(pt).$
@@
    (13) $\mathbf{w}(\tau_{i}`,\sigma`)=\mathbf{w}(\tau_{i},\sigma)-a_{i}\mathbf{w}(\tau,\sigma).$
@@ 4.1.2 Decorated graphs
In this section, we deal with torus fixed points in the moduli stack of stable maps ¯$\overline{\mathcal{M}}_{g,n}(\mathcal{X},\beta)$. Given a stable map $f:(C,x_{1},\cdots,x_{n})\rightarrow\mathcal{X}$ such that $[f:(C,x_{1},\cdots,x_{n})\rightarrow\mathcal{X}]\in\overline{\mathcal{M}}_{g,n}(\mathcal{X},\beta)^{S}$, we will associate with it a decorated graph $\overrightarrow{\Gamma}$. We first give a formal definition.
@@
  • $\sum_{v\in V(\Gamma)}g_{v}+|E(\Gamma)|-|V(\Gamma)|+1=g$.
  • $\sum_{e\in E(\Gamma)}d_{e}[\ell_{\tau_{e}}]=\beta$.
Let $G_{g,n}(\mathcal{X},\beta)$ be the set of all decorated graphs $\overrightarrow{\Gamma}=(\Gamma,\overrightarrow{f},\overrightarrow{d},\overrightarrow{g},\overrightarrow{s})$ that satisfy the above constraints.
@@
the set of edges emanating from $v$ and define $S_{v}=\overrightarrow{s}^{-1}(v)\subset\{1,\ldots,n\}$. The valency of $v$ is defined by $val(v)=|E_{v}|+|S_{v}|$.
Let $f:(C,x_{1},\cdots,x_{n})\in\overline{\mathcal{M}}_{g,n}(\mathcal{X},\beta)^{S}$ be a stable map fixed by the $S$-action. Then for an irreducible component $C^{\prime}\subset C$, $f\rvert_{C^{\prime}}$ either contracts to a point $p_{\sigma}\in\mathcal{X}^{S}$ or $C^{\prime}\cong\mathbb{P}^{1}$ and $f(C^{\prime})=\ell_{\tau}$.
A decorated graph $\overrightarrow{\Gamma}$ associated with $f:(C,x_{1},\ldots,x_{n})\rightarrow\mathcal{X}$ can be defined as follows.
  1. 1.
    We assign a vertex $v$ to each connected component of $f^{-1}(\mathcal{X}^{S})$, and denote the corresponding component by $C_{v}$. Assume that $f(C_{v})=\{p_{\sigma}\}$, then the label map $\overrightarrow{f}(v)=\sigma$, and the genus $\overrightarrow{g}(v)=g_{v}$ is the arithmetic genus of the component $C_{v}$. The marking map $\overrightarrow{s}$ sends the marking points on the component $C_{v}$ to $v\in V(\Gamma)$.
  2. 2.
    We assign an edge $e$ to each connected component $f^{-1}(\mathcal{X}^{1}\backslash\mathcal{X}^{S})$. Let $C_{e}\cong\mathbb{P}^{1}$ be the closure of one connected component in $f^{-1}(\mathcal{X}^{1}\backslash\mathcal{X}^{S})$ and assume that $f(C_{e})=\ell_{\tau}$. We define $\overrightarrow{f}(e)=\tau$.
  3. 3.
@@
The construction above gives a map from ¯$\overline{\mathcal{M}}_{g,n}(\mathcal{X},\beta)^{S}$ to the discrete set $\overrightarrow{\Gamma}\in G_{g,n}(\mathcal{X},\beta)$. Let $\mathcal{F}_{\overrightarrow{\Gamma}}\subset\overline{\mathcal{M}}_{g,n}(\mathcal{X},\beta)^{S}$ denote the preimage of $\overrightarrow{\Gamma}$. Then
@@ 4.1.3 Localization formula on Gromov-Witten theory
where $\Lambda^{\vee}_{g}(u)=\Sigma_{i=0}^{g}(-1)^{i}\lambda_{i}u^{g-i}$ and $\lambda_{i}$ are $\lambda$-classes, see the definition in [liu2011localization, Sec 3.1]. Furthermore, we define
@@
4.2 Torus fixed points
@@
(14) $(Fl)^{S}=\{\vec{I}=(I_{1},\cdots,I_{n})|I_{1}\subset I_{2}\subset\cdots\subset[N_{n+1}],\lvert I_{i}\rvert=N_{i},i=1,\cdots,n\}\,,$
@@
(15) $\lvert(Fl)^{S}\rvert=\prod_{i=1}^{n}\binom{N_{i+1}}{N_{i}}.$
@@
  1. 1.
    If a node $v$ has only one maximal node $w$ smaller than it, the associated matrix of the arrow is non-degenerate and in reduced row/column echelon form whose non-pivots entries vanish.
  2. 2.
@@
Combine the Equation (15) and the Lemma 4.7, and we obtain the following Corollary.
@@
For an arbitrary decorated QP $(\mathbf{Q}=(Q_{0},Q_{1},W),\mathbf{r})\in\Omega^{A}_{n}$, let $\mathcal{Z}$ be the critical locus of $W$. The cardinality of the torus fixed locus is
$\lvert\mathcal{Z}^{S}\rvert=\lvert(Fl)^{S}\rvert=\prod_{i=1}^{n}\binom{N_{i+1}}{N_{i}}.$
According to the Atiyah-Bott localization theorem [ATIYAH19841],
(16) $H_{S}^{*}(\mathcal{Z})\cong\oplus_{p\in\mathcal{Z}^{S}}H^{*}_{S}(p)\,,$
@@
$\phi:H^{*}_{S}(\mathcal{Z})\xrightarrow{\sim}H^{*}_{S}(\mathcal{Z}^{\prime})\,.$
@@
(17) $\gamma\big{|}_{P}=\phi(\gamma)\big{|}_{\varphi(P)},$
@@
It is obvious that the two equivariant cohomology groups are isomorphic by Atiyah-Bott localization theorem (16) and the Corollary 4.8. One can find the mutation $\mu_{v}$ maps $c_{k}^{S}(S_{u}^{\vee})$ to $c_{k}^{S}(S_{u}^{\prime})$ if $\sigma_{u}$ changes its sign according to the Definition 3.4.
@@
(18) $\int_{\beta_{u}}c_{1}(S_{w})=-\operatorname{sign}(\sigma_{w})\delta_{uw}\,.$
Then we have $H_{2}(\mathcal{Z},\mathbb{Z})\cong\mathbb{Z}^{n}$ For two varieties $\mathcal{Z}$ and $\mathcal{Z}^{\prime}$ related by a quiver mutation at the gauge node $v$, we present a transformation of their effective curve classes inspired by (8).
@@
Let $\beta=(\beta_{u})_{u\in Q_{0}\backslash Q_{f}}$ and $\beta^{\prime}=(\beta^{\prime}_{u})_{u\in Q_{0}\backslash Q_{f}}$ be the bases of $H_{2}(\mathcal{Z},\mathbb{Z})$ and $H_{2}(\mathcal{Z}^{\prime},\mathbb{Z})$, respectively, chosen as (18). Then their transformation is given as follows. If $\sigma_{v}>0$, which means $N_{f}(v)>N_{a}(v)$, then
(19) $\phi_{*}(\beta_{u})=\begin{cases}-\beta_{u}^{\prime},&\text{ for }u=v\\ \beta_{u}^{\prime}+\beta_{v}^{\prime}&\text{ for }v\rightarrow u\\ \beta_{u}^{\prime}&\text{ otherwise }\,.\end{cases}$
@@
(20) $\phi_{*}(\beta_{u})=\begin{cases}-\beta_{u}^{\prime},&\text{ for }u=v\\ \beta_{u}^{\prime}+\beta_{v}^{\prime}&\text{ for }u\rightarrow v\\ \beta_{u}^{\prime}&\text{ otherwise }\,.\end{cases}$
@@
Notice that $(c_{1}(S_{u}))_{u\in Q_{0}\backslash Q_{f}}$ spans $H^{2}(\mathcal{Z})$. One can check directly that the $\phi_{*}$ defined above satisfies (8), via Proposition 4.9 and (18). ∎
4.3 One-dimensional torus-fixed orbits
@@ 4.3.1 Grassmannian
    $A=\begin{bmatrix}\vec{v}_{1}&\vec{v}_{2}&\cdots&\vec{v}_{N_{2}}\end{bmatrix}.$
@@
    (21) $\begin{cases}(A_{t})_{ab}=\delta_{f_{a}b}\text{ for }a\neq s\\ (A_{t})_{si_{s}}=\frac{1}{1+t}\\ (A_{t})_{sj_{s}}=\frac{t}{t+1}\end{cases}$
@@
(22) ${Fl(N_{1},\ldots,N_{n})}$${Fl(N_{1},\ldots,N_{n},N_{n+1})}$${Gr(N_{n},N_{n+1})}$$\scriptstyle{\pi}$
@@
  2. 2.
    If $\pi(\ell)$ is a line $\ell_{b}\subset Gr(N_{n},N_{n+1})$, then $\ell_{b}$ is an one-dimensional $S$-orbit closure in $Gr(N_{n},N_{n+1})$ since $\pi$ is $S$-equivariant. By Lemma 4.11, $\ell_{b}$ contains two $S$-fixed points $q_{I},q_{J}\in Gr(N_{n},N_{n+1})^{S}$, satisfying $|I\cap J|=N_{n}-1$. Without loss of generality, we assume that $I\backslash J=\{N_{n+1}-1\}$, $J\backslash I=\{N_{n+1}\}$. Fix a point $q\in\ell_{b}\backslash\{q_{I},q_{J}\}$, and we assume that its matrix representation is given by $A_{t_{0}}$ in (21) for some fixed $t_{0}$. The matrix representation of the fiber $\pi^{-1}(q)$ can be $(A_{1},A_{2},\ldots,A_{n-1},A_{t_{0}})$. Now one can check that there is a $S^{\prime}=(\mathbb{C}^{*})^{N_{n+1}-1}$-action on $\pi^{-1}(q)$, where the inclusion $S^{\prime}\hookrightarrow S$ is given by $(t_{1},t_{2},\ldots,t_{N-2},t,t)$. Actually, this action can be viewed as an extension of the natural $(\mathbb{C}^{*})^{N_{n}}$-action on $Fl(N_{1},N_{2},\ldots,N_{n})$ to $\pi^{-1}(q)$.
    The point $p\in\ell\cap\pi^{-1}(q)$ must be a $S^{\prime}$-fixed point in $Fl(N_{1},N_{2},\ldots,N_{n})$. Otherwise the whole $S$-action on $p$ will result in a larger than one dimensional $S$-orbit, which contradicts the assumption that $\ell$ is one-dimensional. By continuity, when $q$ varies, $p$ should have the same matrix representation. In the formula of (14), either $I_{l}=J_{l}$, or $I_{l}$ contains the $N_{n+1}-1$ and $J_{l}$ contains the $N_{n+1}$. Hence we have proved the induction in case $n$.
@@
    (23) $A_{1}=\begin{bmatrix}s&t&0&\cdots&0&\ldots&0\\ 0&0&1&\cdots&0&\ldots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots&&\vdots\\ 0&0&0&\cdots&1&\ldots&0\end{bmatrix}$
@@
    (24) $A_{2}=\begin{bmatrix}a&0&0&\cdots&0\\ b&0&0&\cdots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ 0&0&1&\cdots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ 0&0&0&\cdots&1\end{bmatrix}.$
@@
We consider a general decorated QP in $\Omega^{A}_{n}$. Let $\mathcal{Z}:=\{dW=0\}\sslash_{\theta}G$ be the associated variety. We denote by $p_{\vec{I}}$ the torus fixed point indexed by a set of sets $\vec{I}=(I_{v})_{v\in Q_{0}\backslash v_{f}}$.
@@
  2. 2.
    There exist at most two maximal nodes that are smaller than $v_{0}$. If there is only one maximal node $v_{1}$ that is smaller than $v_{0}$, then $I_{v_{1}}\cap J_{v_{1}}=r_{v_{1}}-1$. If there are exactly two maximal nodes $v_{1}$ and $v_{2}$ that are smaller than $v_{0}$, then $(I_{v_{k}},J_{v_{k}}),\,k=1,2$ must satisfy one of the following three conditions:
@@
We prove by induction on the number of gauge nodes $n$. For $n=1$ and $2$, this is proved by Lemmas 4.11, 4.12 and 4.13. Assume that the Lemma is true for $n-1$. We’re going to prove the case when the number of gauge node is $n$.
Let $\ell$ be an one-dimensional $S$-orbit closure in $\mathcal{Z}$. When the frozen node $v_{f}$ has only one maximal node $v_{1}$ smaller than it, then we have the following fiber bundle
${\mathcal{Z}^{\prime}}$${\mathcal{Z}}$${Gr(r_{v_{1}},N_{n+1})}$$\scriptstyle{\pi}$
@@
4.4 Behavior of one-dimensional torus-fixed orbits under quiver mutations